Jen invested $7,500 in two back accounts, one paying 4% simple interest, and the other paying 6%. If her annual interest from the two accounts is $420, how much did she invest at 6%?
Answer by mananth(16946)

Part I 6.00% per annum ------------- Amount invested =x
Part II 4.00% per annum ------------ Amount invested = y
7500
Interest----- 420.00

Part I 6.00% per annum ---x
Part II 4.00% per annum ---y
Total investment
x + 1 y= 7500 -------------1
Interest on both investments
6.00% x + 4.00% y= 420
Multiply by 100
6 x + 4 y= 42000.00 --------2
Multiply (1) by -6
we get
-6 x -6 y= -45000.00
Add this to (2)
0 x -2 y= -3000
divide by -2
y = 1500 - x)
Part I 6.00% $ 6000
Part II 4.00% $ 1500

CHECK
6000 --------- 6.00% ------- 360.00
1500 ------------- 4.00% ------- 60.00
Total -------------------- 420.00

m.ananth@hotmail.ca